OpenStudy (anonymous):

Solve: |2n - 3| = 5 A. {-1, 4} B. {1, 4} C. {-1, 1} D. {1, -4

OpenStudy (ajprincess):

|2n - 3| = 5 \(2n-3=\pm5\) 2n-3=5 | 2n-3=-5 solve for n